# On the relationship between Koopman operator approximations and neural ordinary differential equations for data-driven time-evolution predictions

### v1

@@ V Conclusions
Developing and improving upon neural network-based methods for time evolution prediction of dynamical systems is a promising approach for the forecasting of complex dynamical systems. In this work, we have examined the parallels between neural ODE methods and methods based on finite-dimensional approximations of the Koopman operator using neural network-based dictionaries. While methods based on the Koopman operator lead to an approximation of linear dynamics on the space of observables, in practice they often provide poor predictions of the time-evolution of the states due to the difficulty associated with determining a dictionary which spans a Koopman invariant space for a given system. This is especially true for chaotic systems, such as the ones considered here. It has been seen that the state-predictions of these Koopman-based methods can be greatly improved by projecting back to the state space on each timestep. Here we have pointed out that by relifting to the observable space from the state space on each step, we introduce a nonlinearity into the time evolution formulation, resulting in a nonlinear discrete-time map on the state space. Therefore using an EDMD approximation of the operator in this way with a neural network dictionary can be seen as a discrete-time version of a neural ODE. Furthermore, we can also apply a similar projection strategy to the continuous time formulations of EDMD based on the Koopman generator to obtain a continuous time neural ODE directly from the EDMD-DL-optimized operator and dictionary without any further training. Additionally, the model structures arising from EDMD-DL with projection naturally possess a neural network structure that is commonly used in neural ODEs, in which the state is first expanded out to a high-dimensional feature vector (represented by a multi-layer neural network) followed by a linear mapping back to the state space to yield the vector field of the dynamics.
@@
Inspired by these observations, we have implemented several variations of neural ODEs and EDMD-DL, developed by combining different aspects of their respective model structures and training procedures. We provide performance comparisons on two chaotic systems to demonstrate the equivalence of these methods and highlight the aspects of each model that lead to improved performance. Specifically, we show that the novel neural ODE architecture based on the dictionary structure typically used in EDMD-DL, where the state is explicitly separated from the nonlinear dictionary elements, leads to a small improvement in the short-time prediction performance of neural ODEs, as compared to the standard neural ODE architecture. In the context of neural ODEs, separating the linear and nonlinear parts of the vector field provides a nice structure, as the neural network only has to account for the nonlinear dynamics, with the linear terms represented separately. In the examples considered here, the EDMD-structured neural ODE architecture, trained with the usual neural ODE loss, performed best overall, suggesting that this should be the standard structure for neural ODEs.
@@
We see that each of the neural ODE methods and EDMD-DL methods with projection tend to perform comparably on these systems in terms of short-time trajectory prediction, reconstruction of long-time statistics, and the prediction of extreme events; with all of them offering a significant improvement over the standard EDMD-DL linear time evolution on the space of observables. Furthermore, we emphasize that each of these models perform very well in terms of short-time prediction, with less than $3\%$ relative error at $t=\tau_{L}$ for each of the chaotic systems considered here and rare event prediction capability comparable to the non-Markovian approach of Racca and Magri [67]. In addition to accurate short-time tracking, we have also demonstrated that the model predictions are statistically accurate for long times as well, as they yield accurate reconstructions of the long-time probability density functions on the state space and accurately predict the turbulent lifetime for the MFE model of a turbulent shear flow. These findings highlight the promise of neural ODEs and Koopman-based approaches for developing data-driven models capable of generating accurate time evolution predictions in complex and chaotic systems.

### v2

@@ V Conclusions
Developing and improving upon neural network-based methods for time evolution prediction of dynamical systems is a promising approach for the forecasting of complex dynamical systems. In this work, we have examined the parallels between neural ODE methods and methods based on finite-dimensional approximations of the Koopman operator using neural network-based dictionaries. While methods based on the Koopman operator typically lead to an approximation of linear dynamics on the space of observables, in practice these linear methods often provide poor predictions of the time-evolution of the states due to the difficulty associated with determining a dictionary which spans a Koopman invariant space for a given system. This is especially true for chaotic systems, such as the ones considered here. It has been seen that the state-predictions of these Koopman-based methods can be greatly improved by projecting back to the state space on each timestep. Here we have pointed out that by relifting to the observable space from the state space on each step, a nonlinearity is introduced into the EDMD time evolution formulation, resulting in a nonlinear discrete-time map on the state space. Therefore using an EDMD approximation of the operator in this way with a neural network dictionary can be seen as a discrete-time version of a neural ODE. Furthermore, we can also apply a similar projection strategy to the continuous time formulations of EDMD based on the Koopman generator to obtain a continuous time neural ODE directly from the EDMD-DL-optimized operator and dictionary without any further training. Additionally, the model structures arising from EDMD-DL with projection naturally possess a neural network structure that is commonly used in neural ODEs, in which the state is first expanded out to a high-dimensional feature vector (represented by a multi-layer neural network) followed by a linear mapping back to the state space to yield the vector field of the dynamics.
While including the state space projection step in the EDMD-DL timestepping procedure does reduce the model to a neural network representation of the dynamics on the state space, the training procedure and structure of the model are the same as in standard EDMD-DL. Therefore, the model inherits the benefits of the EDMD-DL method. It provides a matrix approximation of the Koopman operator, which can be used to study Koopman eigenvalues, eigenfunctions, and modes [14, 2]. This has been demonstrated in the original EDMD-DL paper [2], where it was also shown that optimizing the basis in EDMD-DL typically yields more accurate and efficient representations of the operator and its spectrum than alternative choices of basis, such as radial basis functions or Hermite polynomials. Many previous works have shown that analysis of Koopman eigenfunctions can be used to glean physical insights into the structure underlying a dynamical system by identifying large scale structures and their associated frequencies in high-dimensional systems, parameterizing basins of attraction, and identifying invariant sets and coherent sets (see, for example 3, 4, 7, 5 for detailed reviews of these applications).
@@
In this work, we have implemented several variations of neural ODEs and EDMD-DL, developed by combining different aspects of their respective model structures and training procedures. We provide performance comparisons on two chaotic systems to demonstrate the equivalence of these methods and highlight the aspects of each model that lead to improved performance. The essentially equivalent performance between EDMD-DL with projection and the neural ODE models, in contrast to the failure of linear EDMD-DL, further emphasizes our main point: the state space projection makes the EDMD-DL model equivalent to a neural ODE in that it gives a nonlinear, neural network representation of the dynamics on the state space. We also show that the novel EDMD-structured neural ODE architecture yields a small improvement in the short-time prediction performance of neural ODEs, as compared to the standard neural ODE architecture. In the context of neural ODEs, separating the linear and nonlinear parts of the vector field provides a nice structure, as the neural network only has to account for the nonlinear dynamics, with the linear terms represented separately. In the examples considered here, the EDMD-structured neural ODE architecture, trained with the usual neural ODE loss, performed best overall, suggesting that this structure can be beneficial to the performance of neural ODEs on chaotic systems such as these.
@@
We see that each of the neural ODE methods and EDMD-DL methods with projection tend to perform comparably on these systems in terms of short-time trajectory prediction, reconstruction of long-time statistics, and the prediction of extreme events; with all of them offering a significant improvement over the standard EDMD-DL linear time evolution on the space of observables. Furthermore, each of these models performs very well in terms of short-time prediction, with less than $3\%$ relative error at $t=\tau_{L}$ for each of the chaotic systems considered here and rare event prediction capability comparable to the non-Markovian approach of Racca and Magri [51]. In addition to accurate short-time tracking, we have also demonstrated that the model predictions are statistically accurate for long times as well, as they yield accurate reconstructions of the long-time probability density functions on the state space and accurately predict the turbulent lifetime for the MFE model of a turbulent shear flow. These findings highlight the promise of neural ODEs and Koopman-based approaches for developing data-driven models capable of generating accurate time evolution predictions in complex and chaotic systems.